# SpikySpace: A Spiking State Space Model for Energy-Efficient Time Series Forecasting

### v1

@@ I Introduction
  • We present SpikySpace, a hardware-friendly model for time series tasks implemented using a fully spiking state model. To our knowledge, SpikySpace is the first fully spiking SSM, where both the latent state updates and nonlinear transformations are realized through discrete spike-driven computation. This makes it exceptionally energy-efficient and inherently compatible with neuromorphic hardware.
  • We propose lightweight approximations for SiLU and Softplus, enabling full compatibility with neuromorphic hardware, removing a major obstacle to deploying SSM-based SNNs on energy constraint systems.
  • SpikySpace achieves competitive or superior forecasting accuracy while drastically reducing energy consumption and parameter size compare with the state-of-the-art methods, demonstrating a highly favorable balance between accuracy and efficiency for neuromorphic and edge deployment.
II Related Work
II-A Spiking Neural Networks
Spiking neural networks (SNNs) consist of interconnected spiking neurons that communicate through discrete spike events, enabling energy-efficient and event-driven computation [12, 13]. A spiking neuron integrates incoming signals over time and emits a spike once its “membrane potential” exceeds a threshold, mimicking the behavior of biological neurons. Because information is represented as discrete spike events rather than continuous activations, computation only occurs when events happen, leading to lower energy consumption and improved efficiency. Moreover, owing to their temporal coding mechanism, SNNs are particularly well-suited for modeling temporal or sequential signals. Such network architectures have been extensively explored across various tasks, ranging from image classification and natural language processing to autonomous control [14, 15, 16, 17, 18].
Although many SNN architectures employ transformer-style self-attention blocks to exploit long-range dependencies, the underlying attention mechanism still imposes a quadratic complexity barrier in sequence length. To address this limitation, a growing line of research has explored State-Space Models (SSMs) such as S4 and Mamba as efficient alternatives to attention. These models achieve linear-time (or near-linear) complexity and enable fully parallel updates with respect to input length [4, 5].
To capture long-range dependencies, many SNN architectures often utilise transformer-like self-attention mechanisms. However, the attention operation involves quadratic computational and memory complexity with respect to sequence length, and the dense all-to-all interactions are not well aligned with the sparse and event-driven nature of SNNs, causing suboptimal efficiency. Alternative architectures such as State Space Models (SSMs), including S4 and Mamba [4, 5], provide linear-time sequence modeling and parallel updates, but adapting these continuous-time formulations to the spiking domain remains non-trivial challenges.
II-B Challenges of Adapting SSMs to SNNs
Structured SSM architectures are originally designed for conventional artificial neural networks (ANNs). Their architectures and operations are not directly compatible with SNNs. SSM-based layers typically rely on real-valued state vectors that evolve at every discrete time step or through continuous differential equations and use dense linear algebra operations such as matrix multiplications and convolutions. In contrast, SNNs operate in an event-driven manner with binary spikes and sparse, asynchronous updates. The continuous-time dynamics and long convolutions of SSMs do not naturally align with this spike-driven computation model.
There have been some explorations of integrating SNNs with SSMs. Existing studies [6, 7, 8, 9] have made progress toward spiking variants of SSMs. However, their formulations still rely on continuous states and dense multiplication operations, which prevent them from fully exploiting the sparsity and event-driven benefits of spiking computation. Additionally, many of these designs treat SNNs merely as wrappers or feature extractors around conventional SSM blocks. Complex activation functions and operations used in SSMs, such as SoftPlus and SiLU, are retained in a continuous manner.
Overall, current solutions fail to achieve a genuinely end-to-end, fully spiking state-space model in which the latent state is inherently updated by sparse spike activity.
In SpikySpace, we address this gap by exploring how the principled dynamics of SSMs can be adapted to the spiking regime, paving the way for a Spiking SSM architecture that combines the long-range modeling capability of SSMs with the efficiency of event-driven SNN computation.
II-C Time-series Forecasting
Time-series forecasting aims to predict future values based on historical sequences. It serves as a fundamental technique across diverse domains, including finance, healthcare, and industrial monitoring.
Classical statistical methods such as ARIMA [19] and exponential smoothing [20] capture trend and seasonality under stationarity assumptions, but struggle with nonlinear dynamics, regime shifts, and long-range dependencies in multivariate settings.
Neural networks such as LSTMs [21] and GRUs [22] improve the modeling of sequential dependencies but can be computationally intensive for long sequences. Temporal convolutional networks (TCNs) [23] enlarge receptive fields through dilation and allow parallel training, but require deep stacks or large kernels to capture long-range contexts. Transformers have also been applied to time-series forecasting. Variants such as Informer [24] and Autoformer [25] reducing computation through sparse attention or decomposition strategies and show strong accuracy on long-horizon benchmarks [26]. However, these models remain memory-bound due to key–value caches and dense step-wise updates.
Complementary efficiency techniques for neural networks, such as quantization, pruning, lightweight backbones, and mixed precision, can reduce arithmetic cost and memory footprint. However, the dense computation pattern remains a fundamental limitation: compute and memory traffic still scale with sequence length rather than information content [27]. As a result, throughput and latency remain sensitive to context windows and key–value caches even after compression, particularly in streaming or on-device settings.
To enable truly efficient forecasting, recent studies have explores SNNs for sequential and multivariate prediction, leveraging event-driven sparsity to align computation with informative temporal changes  [2, 28, 3, 29]. Recurrent and convolutional spiking models capture local dynamics with sparse updates, while spiking Transformers introduce positional or relative attention to model longer-range dependencies. Graph-based spiking architectures have also been proposed to exploit spatial correlations in sensor networks [28, 29].
However, these approaches often retain per-timestep (or per-node) dense operations at some stage of the pipeline and trade accuracy for sparsity on long horizons. This motivates our direction: combining the structured memory of state-space models with event-triggered execution, so that updates are driven by spike activity rather than timestamp, preserving long-range accuracy while achieving sparse, and information-aligned computation.
III Preliminary
III-A Problem Statement
Definition 1 (Multivariate Time Series): A multivariate time series is an ordered sequence of observations $X=\{x_{1},x_{2},\ldots,x_{|X|}\}\in\mathbb{R}^{|X|\times N}$, where $N$ is the number of variables and each observation $x_{i}\in\mathbb{R}^{N}$ represents the values of all variables at time step $i$. Given a historical window $X_{H}\in\mathbb{R}^{H\times N}$, the goal is to forecast the next $G$ steps. The target sequence is $X_{G}\in\mathbb{R}^{G\times N}$ that represents the future states of all variables. During training, $(X_{H},X_{G})$ pairs are sampled from historical data for the model to learn temporal dependencies. During inference, an observed series $X_{O}\in\mathbb{R}^{O\times N}$ is provided, and the model produces predictions for the next $G$ steps. In this work, the mapping $X_{H}\mapsto X_{G}$ is implemented by an SNN, which performs event-driven updates for efficient computation on streaming time-series data.
III-B Spike Neuron Model
SNNs communicate via discrete spikes rather than continuous activation, leading to event-driven and temporally sparse computation. However, training SNNs is non-trivial because the spike generation function is inherently non-differentiable, which makes gradient-based optimization inapplicable. To overcome this, two main strategies are commonly employed. The first relies on surrogate gradients [30], which approximate the gradients of spike events to allow backpropagation. The second, ANN-to-SNN conversion [31], leverages a pre-trained artificial neural network (ANN) and maps its continuous neuron activations to spiking dynamics, producing a functionally similar SNN. In this work, we follow the ANN-to-SNN conversion approach to construct our SpikySpace.
Among spiking neuron models, the Integrate-and-Fire (IF) neuron is widely used for its simplicity and efficiency [32]. In the IF model, a neuron’s membrane potential $V(t)$ integrates incoming synaptic currents over time:
$\frac{dV(t)}{dt}=I_{\text{syn}}(t),$ (1)
[…]

### v2

@@ 1 Introduction
  • We propose SpikySpace, a novel architecture that integrates the continuous dynamics of SSMs into an event-driven framework for time series tasks. By implementing a spike-driven selective scan constrained to powers of two, we achieve a recurrent backbone with linear complexity that is completely free of multiplication.
  • We design and theoretically validate two hardware-friendly approximations, PTsoftplus and PTSiLU. Constructed from linear transformations and power-of-two components, these operators are proven to approximate the original functions closely while eliminating costly exponentials and divisions, removing a major obstacle to deploying SSM-based SNNs on energy-constrained systems.
  • We validated SpikySpace on multivariate forecasting benchmarks, and show that it achieves competitive accuracy with substantial savings in energy and parameter size. We believe this is evidence that SpikySpace is a highly practical solution for edge deployment.
2 Related Work
Spiking Neural Networks for Sequence Modeling.
Spiking Neural Networks (SNNs) offer significant energy efficiency through event-driven computation (Taherkhani et al., 2020; Zhou et al., 2024). While successful in tasks ranging from image classification (Hu et al., 2023; Tang et al., 2024a) to NLP (Zhou et al., 2023; Tang et al., 2024b) with SNNs’ intrinsic temporal dynamics, modeling long sequences remains a challenge. Current Spiking Transformers capture global dependencies but suffer from the quadratic complexity of attention mechanisms, which undermines the sparse efficiency of neuromorphic computing. This bottleneck necessitates exploring linear-complexity backbones, such as State Space Models (SSMs) (Gu et al., 2021), to better align with the efficiency goals of SNNs.
Spiking State Space Models.
Adapting the continuous, dense dynamics of SSMs to the discrete, sparse domain of SNNs presents significant challenges. Although recent studies have explored integrating SNNs with SSMs (Shen et al., 2025; Bal and Sengupta, 2024; Zhong et al., 2024; Huang et al., 2025), these implementations are predominantly hybrid. Most retain real-valued state updates, dense matrix multiplications, or complex continuous activations like Softplus and SiLU within the recurrent loop. Such designs fail to fully exploit the energy benefits of neuromorphic hardware, as the core state evolution remains reliant on floating-point arithmetic. In contrast, SpikySpace proposes a fully spiking backbone where latent states are updated via sparse bit shift operations and additions.
Time-series Forecasting.
Forecasting time series is critical for domains ranging from industrial monitoring to energy management. While transformers (Zhou et al., 2021; Wu et al., 2022; Nie et al., 2023) currently forecast with high accuracy, their dense attention maps and KV-cache requirements make them memory-bound and prohibitive for deployment on the edge. Model compression techniques like quantization and pruning (Cheng et al., 2020) alleviate some overhead but do not alter the underlying dense computation. Consequently, SNNs have been investigated for efficient forecasting (Lv et al., 2024; SHIBO et al., 2025; Hu et al., 2025). However, existing SNN forecasters often face a trade-off where they either rely on structures similar to RNNs that struggle with long horizons or adopt transformer architectures that suffer from quadratic complexity. There remains a lack of neuromorphic forecasting solutions that are both highly accurate and energy efficiency.
3 Preliminary
Problem Statement.
Let $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{T}\}\in\mathbb{R}^{T\times N}$ denote a multivariate time series with $N$ variables. Given a historical window of length $L$, $\mathbf{X}_{in}\in\mathbb{R}^{L\times N}$, the goal is to forecast the future horizon $H$, denoted as $\mathbf{Y}_{pred}\in\mathbb{R}^{H\times N}$. We aim to learn a mapping $f_{\theta}:\mathbf{X}_{in}\mapsto\mathbf{Y}_{pred}$ parameterized by an SNN, leveraging event-driven computation to process the temporal stream efficiently.
Spiking Neural Networks.
SNNs process information through discrete spike events $s_{t}\in\{0,1\}$ rather than continuous activations. To bypass the training difficulties caused by non-differentiable spike generation, we utilize the ANN-to-SNN conversion framework, which maps the weight of a pre-trained ANN to a spiking equivalent. Our architecture is built upon the Integrate-and-Fire (IF) neuron. In the discrete time domain, the membrane potential $V_{t}$ simply accumulates synaptic input $I_{t}$. Upon exceeding a threshold $\theta$, the neuron fires as $s_{t}=1$ and resets via subtraction as $V_{t}\leftarrow V_{t}-\theta$. For hardware efficiency, we adopt the Average Integrate-and-Fire (Avg-IF) variant (Yan et al., 2025), which optimizes memory access patterns by synchronizing potential updates. Detailed algorithms are provided in Appendix A. In this work, we denote the spike neuron used to generate spike trains as $\mathcal{SN}(\cdot)$, which accumulates synaptic inputs and triggers a binary output upon exceeding a firing threshold $\theta$.
State Space Models.
State Space Models (SSMs) capture sequence dynamics through a latent state $h_{t}\in\mathbb{R}^{D}$. While theoretically grounded in continuous differential equations, efficient inference relies on the discretized recurrence:
$h_{t}=\bar{\mathbf{A}}h_{t-1}+\bar{\mathbf{B}}x_{t},\quad y_{t}=\mathbf{C}h_{t},$ (1)
where $\bar{\mathbf{A}}$ and $\bar{\mathbf{B}}$ are the discretized system parameters. We provide the detailed derivation from the continuous differential equations to this discrete form in Appendix B. As Eq. (1) relies on dense floating-point multiplications for both state transitions and input projections. SpikySpace eliminates these bottlenecks through a strategic combination of spike-based encoding and power-of-two quantization.
4 Method
4.1 Overall Framework
Figure 1: The overall architecture of SpikySpace. (Left) The model consists of stacked Spiking Mamba blocks followed by a prediction head. Each block employs a standard pre-norm residual design. Normalization is performed via RMSNorm, which admits hardware-efficient approximations (Tang et al., 2024b) for deployment. (Right) Inside the block, input features are projected and split into two branches comprising a residual gating branch and a data branch processed by our Spiking SSM. Weights are quantized to INT8, and activations are encoded as spikes, enabling an inference pipeline free of multiplication.
As illustrated in Figure 1, SpikySpace stacks multiple Spiking Mamba blocks to process multivariate time series. Each block integrates our proposed spiking scan mechanism within a residual backbone to capture temporal dependencies. Finally, a linear head maps the learned representations to the target prediction horizon.
We implement a co-design strategy focused on efficiency. Specifically, SSMs weight are quantized to 8-bit integers via post-training quantization, while dynamic activations are encoded as binary spikes. This combination allows us to replace dense matrix products with sparse accumulations of integer weights of lower bitwidths. Furthermore, we replace expensive operations in standard activations such as Softplus and SiLU with approximations based on bit shifts, resulting in a purely additive SSMs inference ideally suited for energy-constrained edge devices.
4.2 State Space Module
The Spiking State Space Module serves as the core component of SpikySpace, adapting the continuous dynamics of Mamba into a sparse mechanism driven by events. Functionally, we organize this module into sequential stages comprising convolutional spike encoding, selective parameterization, the spiking selective scan, and output gating.
Local Feature Encoding.
Given the input data path $\mathbf{x}_{\text{in}}\in\mathbb{R}^{B\times L\times D}$, we first encode the continuous features into an initial spike train $\mathbf{s}_{\text{in}}$ via the spike neuron $\mathcal{SN}$. Subsequently, to capture local temporal contexts, $\mathbf{s}_{\text{in}}$ undergoes a depth-wise separable 1D convolution followed by another spike neuron. The encoding process is formalized as:
$\mathbf{s_{\text{in}}}=\mathcal{SN}(\mathbf{x}_{\text{in}}),$ (2)
$\mathbf{s}=\mathcal{SN}(\text{Conv1D}(\mathbf{s}_{\text{in}}),$ (3)
Here, $\mathbf{s}\in\{0,1\}^{B\times L\times D}$ serves as the sparse input to the subsequent SSM.
Selective Parameterization.
Given the input spike train $\mathbf{s}$, we instantiate a latent dynamical system governed by the parameters $(\boldsymbol{\Delta},\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$. Specifically, $\boldsymbol{\Delta}$ represents the discretization step size, while $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ define the state evolution and linear projections.
Let $N$ be the SSM state dimension and $r_{\Delta}$ be the rank of the step-size projection. We first apply a linear transformation with weights $\mathbf{W}\in\mathbb{R}^{d\times(r_{\Delta}+2N)}$ to the input $\mathbf{s}$ and split the result into the projection weights $\mathbf{B},\mathbf{C}\in\mathbb{R}^{B\times L\times N}$ and the raw step size $\boldsymbol{\Delta}_{\text{raw}}\in\mathbb{R}^{B\times L\times r_{\Delta}}$. Then $\boldsymbol{\Delta}_{\text{raw}}$ is passed through a spiking neuron to generate intermediate spikes, denoted as $\boldsymbol{\Delta}_{\text{mid}}$. These spikes are then processed via a quantized linear transformation and fed into our proposed PTSoftplus. Finally, a second spiking neuron encodes the output into the binary gating signal $\boldsymbol{\Delta}_{s}$. The entire parameter generation process is formulated as:
$\boldsymbol{\Delta}_{\text{raw}},\mathbf{B},\mathbf{C}$ $=\text{Split}(\mathbf{W}\mathbf{s}+\mathbf{b}),\quad\boldsymbol{\Delta}_{\text{mid}}=\mathcal{SN}\!\left(\boldsymbol{\Delta}_{\text{raw}}\right),$ (4)
$\tilde{\boldsymbol{\Delta}}$ $=\mathcal{Q}(\mathbf{W}_{\Delta}\boldsymbol{\Delta}_{\text{mid}}+\mathbf{b}_{\Delta}),$
$\boldsymbol{\Delta}_{\text{out}}$ $=\mathrm{PTSoftplus}\!\left(\tilde{\boldsymbol{\Delta}}\right),\quad\boldsymbol{\Delta}_{s}=\mathcal{SN}\!\left(\boldsymbol{\Delta}_{\text{out}}\right),$
Here, $\mathrm{PTSoftplus}$ is our hardware-friendly approximation of $\mathrm{Softplus}$ introduced in Section 4.3, and $\mathcal{Q}$ represents the quantization process as Equation 13. Together with the learnable parameters $\mathbf{A}$ and $\mathbf{D}$, these inputs are forwarded to the selective scan module.
Spiking Selective Scan.
With the spiking step size $\boldsymbol{\Delta}_{s}$ and discrete parameters, the system enters the recursive scan stage. In standard Mamba, the transition matrix is obtained through an exponential operation $\overline{\mathbf{A}}_{t}=\exp(\boldsymbol{\Delta}_{t}\mathbf{A})$, which requires computationally intensive floating-point exponentiation. In SpikySpace, the discretization is simplified by the binary nature of the step size $\boldsymbol{\Delta}_{s}$. Instead of continuous decay, the state update becomes a conditional bit-shift operation applied to the hidden state. Specifically, we pre-compute the shift amount $\mathbf{K}=\lfloor\mathbf{A}\rceil$ derived from the static dynamics parameter. At each time step, if the spike $\boldsymbol{\Delta}_{s}$ is active as a $1$, the state undergoes a bit-shift; otherwise, it remains preserved. This effectively replaces the computationally expensive exponential operation with hardware-intrinsic bit operations:
$\overline{\mathbf{A}}_{t}=\begin{cases}2^{\mathbf{K}}&\text{if }\boldsymbol{\Delta}_{s,t}=1,\\ \mathbf{I}&\text{if }\boldsymbol{\Delta}_{s,t}=0.\end{cases}$ (5)
Importantly, replacing the natural exponential with a base-2 formulation does not reduce the expressive power of the model. Since $2^{x}=e^{\ln(2)x}$, the above transition is equivalent to scaling the effective step size by a constant factor $\ln(2)$. This corresponds to a fixed rescaling of the temporal dynamics, which can be fully compensated by the learnable parameters during training.
To maintain sparsity, we enforce the hidden state $\mathbf{h}_{t}$ to be event-driven. The recurrent dynamics at timestep $t$ are defined as:
$\overline{\mathbf{B}}_{t}$ $=\boldsymbol{\Delta}_{s,t}\mathbf{B}_{t},$ (6)
$\mathbf{h}_{t}$ $=\mathcal{SN}(\overline{\mathbf{A}}_{t}\mathbf{h}_{t-1}+\overline{\mathbf{B}}_{t}\mathbf{s}_{t}),$ (7)
Here, $\mathbf{h}_{t}\in\{0,1\}^{B\times d_{\text{hidden}}\times n\times T}$ denotes the binary hidden state. Since $\mathbf{s}_{t}$ is binary, the input term $\overline{\mathbf{B}}_{t}\mathbf{s}_{t}$ reduces to sparse accumulation.
Finally, the module output $\mathbf{y}_{t}$ is computed via the dynamic projection $\mathbf{C}_{t}$ and the skip connection $\mathbf{D}$:
This entire scanning mechanism achieves $O(L)$ complexity while operating strictly with accumulations and bit-shifts, making it highly efficient for edge deployment. The detailed procedure is summarized in Algorithm 1.
Algorithm 1 SpikingMamba with Selective Scan
1:Input: Spike train $\mathbf{s}\in\{0,1\}^{B\times L\times d}$; sequence length $L$
2:Require: Linear Projection Weight $\mathbf{W}$, $\mathbf{W}_{\Delta}$, Bias $\mathbf{b}$, $\mathbf{b}_{\Delta}$, SSM dynamics $\mathbf{A}\leftarrow-\exp(\mathbf{A}_{\log})$,  $\mathbf{D}$
3:Output: Predicted Spike Train $\mathbf{y}=\{\mathbf{y}_{t}\}_{t=1}^{L}$
4:Stage 1: Parameters Preparation
[…]
11:Stage 2: Selective Scan
12:// Keep $\boldsymbol{\Delta}_{s},h_{t},s_{t}$ as spikes
13:for $t=1$ toLL do
14:  $\overline{\mathbf{A}}_{t}=\begin{cases}2^{\mathbf{K}},\mathbf{K}=\lfloor A\rceil&\text{if }\boldsymbol{\Delta}_{s,t}=1,\\ \mathbf{I}&\text{if }\boldsymbol{\Delta}_{s,t}=0.\end{cases}$
15:  $\overline{\mathbf{B}}_{t}\leftarrow\begin{cases}\mathbf{B}_{t}&\text{if }\boldsymbol{\Delta}_{s,t}=1,\\ 0&\text{if }\boldsymbol{\Delta}_{s,t}=0.\end{cases}$
16:  $\mathbf{h}_{t}\leftarrow\mathcal{SN}(\overline{\mathbf{A}}_{t}\mathbf{h}_{t-1}+\overline{\mathbf{B}}_{t}\mathbf{s}_{t})$
18:end for
19:return $\mathbf{y}$
Output Gating.
The final stage employs a residual branch to stabilize the learned representations. The residual signal $\mathbf{x}_{\text{res}}$ is processed through quantization and our PTSiLU activation, before being modulated by the scan output $\mathbf{y}$, followed by an output linear projection:
$\mathbf{\tilde{y}}=\mathbf{y}\odot\text{PTSiLU}(\mathcal{Q}\!\left(\mathbf{x}_{\text{res}}\right))$ (9)
$\mathbf{z}=\mathbf{W}_{\text{o}}\mathbf{\tilde{y}}+\mathbf{b}_{\text{o}}.$ (10)
Here, the binary spike train $\mathbf{y}$ acts as a sparse mask. As introduced in Section 4.4, PTSiLU employs power-of-two scaling, the non-zero elements in $\tilde{\mathbf{y}}$ remain compatible with bit-wise logic. This transforms the final projection with the weight of $\mathbf{W}{\text{o}}$ into a sparse shift-and-accumulate operation, allowing the entire block to bypass dense multiplication.
4.3 Power-of-Two Softplus
Standard SSMs rely on the Softplus function, defined as $\mathrm{Softplus}(x)=\ln(1+e^{x})$, to ensure the positivity of the timescale parameter $\boldsymbol{\Delta}$. As Softplus depends heavily on operations like exponential and logarithmic that are computationally prohibitive for neuromorphic hardware, we propose Power-of-Two Softplus(PTSoftplus) to resolve this. By utilizing exclusively power-of-two scaling, PTSoftplus replaces expensive floating-point units with efficient bit-shifts and additions, while preserving the convexity required for stable dynamics. Specifically, it is defined as the following function:
[…]
To guarantee $C^{1}$ continuity at the junction $x_{c}$, the constants are analytically derived as:
Theoretical Properties.
We provide theoretical guarantees to ensure that our approximation supports stable gradient-based training.
Lemma 1.
PTSoftplus is continuously differentiable.
The proof of Lemma 1 is deferred to Appendix C.1. Next, we demonstrate that the proposed PTSoftplus function closely approximates Softplus. Specifically, the functional deviation is bounded by a small constant. The derivatives of PTSoftplus and Softplus also exhibit pointwise closeness, ensuring the approximation does not detrimentally affect the model’s training dynamics.
Lemma 2.
$\|{\rm PTSoftplus}-{\rm Softplus}\|_{\infty}\leq 0.914.$
$\|{\rm PTSoftplus}^{\prime}-{\rm Softplus}^{\prime}\|_{\infty}\leq 0.371.$
The detailed proof of Lemma 2 is relegated to Appendix C.2. Fig. 2 visually demonstrates the function shapes and the deviations between the two functions and their respective derivatives.
4.4 Power-of-Two SiLU
Similarly, the output gating branch relies on the SiLU activation, defined as $\mathrm{SiLU}(x)=x\cdot\sigma(x)$, sharing similar hardware bottlenecks as Softplus. Thus, we propose Power-of-Two SiLU (PTSiLU), a piecewise approximation tailored for hardware consistency defined as the following function:
where the parameters
Theoretical Properties.
Similar to PTSoftplus, we first provide theoretical guarantees to ensure that PTSiLU supports stable gradient-based training.
Lemma 3.
PTSiLU is continuously differentiable.
The proof of Lemma 3 is deferred to Appendix C.3. Next, we show that PTSiLU closely approximates SiLU, evidenced by the bounded difference between their function values. Furthermore, the proximity of their first-order derivatives ensures that this approximation does not impair the convergence or stability of the model’s training process.
Lemma 4.
$\|{\rm PTSiLU}-{\rm SiLU}\|_{\infty}\leq 0.316.$
$\|{\rm PTSiLU}^{\prime}-{\rm SiLU}^{\prime}\|_{\infty}\leq 0.263.$
The proof of Lemma 4 is in Appendix C.4. Fig. 3 visually confirms the small deviations in both the function shapes and the derivatives of the two functions.
4.5 Training Process
As we adopted an ANN-to-SNN conversion method to obtain our SpikySpace, we first trained a quantized ANN model with PTsoftplus and PTSiLU as the activation functions and then converted it into an SNN model. For activation quantization, we follow BiT (Liu et al., 2022) and Sorbet (Tang et al., 2024b) to employ Learned Step-size Quantization (LSQ). The quantized activation $X_{Q}$ is derived from the real-valued input $X_{R}$ as:
$\mathcal{Q}(X_{R})=\alpha\cdot\operatorname{clip}\!\left(\operatorname{round}\!\left(\frac{X_{R}-\beta}{\alpha}\right),Q_{n},\;Q_{p}\right)+\beta$ (13)
where $\alpha$ is the learnable step size and $\beta$ is the offset which can be further quantized to power-of-two values without accuracy loss as discussed in (Tang et al., 2024b). Detailed settings are provided in Appendix D.5.
Table 1: Experimental results of time-series forecasting across four datasets. Models are listed as rows. For each dataset, results include horizons $L\in\{3,6,12,24\}$. The best and second-best results in SNNs are highlighted in bold and underlined. $\uparrow$’ means ‘the higher the better’, and $\downarrow$’ means the opposite, i.e., ‘the lower the better’.
Model SNN Metric Metr-la Pems-bay Solar Electricity Avg.
3 6 12 24 3 6 12 24 3 6 12 24 3 6 12 24
GRU $\times$ $R^{2}\uparrow$ .803 .761 .682 .614 .783 .769 .696 .696 .962 .950 .907 .875 .983 .981 .980 .972 .834
[…]
iTransformer $\times$ $R^{2}\uparrow$ .864 .849 .763 .538 .938 .888 .797 .629 .974 .964 .918 .879 .983 .977 .977 .977 .870
[…]
Mamba $\times$ $R^{2}\uparrow$ .891 .842 .754 .612 .922 .875 .778 .606 .973 .954 .915 .824 .989 .987 .986 .984 .868
[…]
SpikeTCN $\checkmark$ $R^{2}\uparrow$ .845 .799 .718 .602 .862 .829 .782 .681 .946 .937 .893 .840 .974 .970 .968 .963 .851
[…]
SpikeRNN $\checkmark$ $R^{2}\uparrow$ .784 .731 .661 .557 .763 .721 .710 .693 .933 .923 .903 .820 .984 .978 .979 .964 .819
[…]
iSpikformer $\checkmark$ $R^{2}\uparrow$ .805 .765 .723 .549 .935 .884 .787 .622 .972 .955 .918 .869 .982 .974 .973 .974 .856
[…]
SpikeSTAG $\checkmark$ $R^{2}\uparrow$ .873 .822 .734 .590 .874 .835 .787 .626 .973 .950 .926 .879 .987 .986 .985 .984 .863
[…]
SpikySpace $\checkmark$ $R^{2}\uparrow$ .895 .847 .760 .610 .921 .866 .761 .605 .971 .953 .914 .825 .994 .992 .990 .990 .869
[…]
5 Result
We evaluate SpikySpace on four multivariate time-series benchmarks across traffic, solar energy, and electricity domains. The comparison includes mainstream ANNs such as GRU and iTransformer, alongside state-of-the-art SNNs like iSpikformer and SpikeSTAG. We report the Coefficient of Determination $R^{2}$ and Root Relative Squared Error (RRSE) across prediction horizons of {3, 6, 12, 24} steps following previous works. All experiments utilize an NVIDIA A100 GPU with a timestep of $T=3$. Dataset specifications and hyperparameter details are provided in Appendix D.
5.1 Comparing with the Baseline
As shown in Table 1, SpikySpace achieves state-of-the-art SNN performance, ranking first on two datasets and remaining highly competitive on the others. On the Electricity dataset, our model achieves an average $R^{2}$ of 0.992, even outperforming dense iTransformer at 0.978. This result indicates that our method effectively preserves high-precision dynamics despite the quantization. On traffic dataset Metr-la, SpikySpace surpasses the strongest SNN baseline by a margin of 3.0% on average (0.778 vs 0.755). This confirms that the spiking selective scan can capture long-range dependencies better than traditional recurrence. These results demonstrate that SpikySpace closes the accuracy gap between SNNs and dense ANNs while maintaining neuromorphic efficiency.
5.2 Energy Analysis
Table 2: Comparison of the model size, energy consumption, and performance across SpikySpace and baselines. The energy reduction is calculated relative to that of iTransformer. The measurements are based on the Electricity dataset with a prediction horizon of 3.
SpikySpace SSM 0.868 0.17 98.20% 0.994 0.137
SpikeSTAG LSTM, transformer 1.566 4.39 53.64% 0.987 0.207
iSpikformer transformer 1.634 3.19 66.30% 0.982 0.214
iTransformer transformer $\times$ 1.634 9.47 / 0.983 0.213
The computational cost of SNNs depends on the time-window length $T$ and the spike rate $s$ (Yan et al., 2024). Unlike prior SNN works that estimate energy solely based on synaptic operations, our evaluation explicitly accounts for the hardware bottlenecks, such as memory access and state updates. Specifically, the energy of a spiking layer can be represented as:
$E_{\text{total}}=T\cdot s\cdot E_{\text{ACC}}+T\cdot s\cdot(E_{\text{move}}+E_{\text{weight}})$ (14)
, where $E_{\text{ACC}}$ is the accumulation energy, $E_{\text{move}}$ is the spike movement energy and $E_{\text{weight}}$ is the SRAM weight access energy. All the results are based on the energy data we measured on a commercial 22nm process.
However, in our SpikySpace, the recursive accumulation of the state transition $\overline{\mathbf{A}}_{t}$ introduces a positive value and tends to saturate the membrane potential. Without intervention, this drives the firing rate toward $\sim\!100\%$, where spikes occur at every timestep, negating the sparsity benefit. To address this, we implement a temporal rescaling strategy. The firing threshold is scaled $T$ times with weights adjusted correspondingly to be $T$ times larger. This effectively aggregates the information of $T$ steps, reducing the firing frequency from $1$ to $1/T$ while maintaining the consistency of the output signal magnitude. Empirically, this strategy drastically reduces the average spike rate from $66.0\%$ to $26.8\%$.
As shown in Table 2, With a model size that is only 53.1% to 55.4% of the baselines, SpikySpace reduced estimated energy consumption by 94.7%-98.2%, consuming only 0.17 mJ on the Electricity dataset.
5.3 Ablation Study
We will now study the effectiveness of our activation approximations and the model’s sensitivity to timestep variations.
Table 3: Ablation study results of PTSoftplus and PTSiLU on full-precision ANNs.
Model Variant Metric Prediction Horizon
3 6 12 24 Avg.
SSM-softplus-SiLU $R^{2}\uparrow$ .895 .846 .763 .617 .780
RRSE $\downarrow$ .342 .414 .514 .653 .481
SSM-PTsoftplus-SiLU $R^{2}\uparrow$ .892 .845 .760 .619 .779
RRSE $\downarrow$ .347 .416 .517 .651 .483
SSM-softplus-PTSiLU $R^{2}\uparrow$ .895 .848 .752 .613 .777
RRSE $\downarrow$ .342 .412 .525 .656 .484
SSM-PTsoftmax-PTSiLU $R^{2}\uparrow$ .893 .847 .755 .614 .777
RRSE $\downarrow$ .346 .413 .522 .656 .484
Table 4: Ablation study on the impact of PTsoftplus and PTSiLU. ‘Bits’ in this table denotes the activation quantization bit-width.
Bits PTsoftplus PTSiLU $R^{2}$ $\Delta R^{2}$
4 $\times$ $\times$ .607 N/A
$\checkmark$ $\times$ .611 +.004
$\times$ $\checkmark$ .611 +.004
$\checkmark$ $\checkmark$ .610 +.003
1 $\times$ $\times$ .612 N/A
$\checkmark$ $\times$ .605 -.007
$\times$ $\checkmark$ .616 +.004
$\checkmark$ $\checkmark$ .603 -.009
Approximated Operations. Table 3 shows that replacing standard Softplus and SiLU with PTSoftplus and PTSiLU causes negligible performance drops across all horizons, confirming the effectiveness of our activation approximation. We further quantify the impact under different activation bit-widths and demonstrate that our variants remain stable even under extremely low-bit settings, making them highly suitable for efficient hardware implementation. The results are shown in Table 4.
Timestep. We then analyze the influence of the simulation timestep $T$. As shown in Figure 4, the performance is stable as $T$ varies from 1 to 15. The curves for both Solar and Metr-la datasets are nearly flat, indicating that the model is robust to different update granularities. We adopt $T=3$ as the default setting to balance computational cost and stability at long prediction horizons.
Figure 4: Ablation study results of different timestep parameters across horizons on Solar and Metr-la datasets.
6 Discussion
Neuromorphic Compatibility. To assess hardware feasibility, we implemented the core operators within the Lava framework targeting Intel’s Loihi. Going beyond standard theoretical estimates, we conducted a rigorous energy evaluation of our operators on a 22nm process. Unlike prior works that focus solely on computational costs, our analysis explicitly accounts for data movement overheads. By incorporating these memory access dynamics, we provide a fine-grained and more realistic energy profile, ensuring that our reported efficiency gains translate effectively to physical hardware implementations. Future work will focus on deploying SpikySpace on physical platforms like Loihi (Davies et al., 2018) or BrainScaleS-2 (Pehle et al., 2022) to obtain empirical latency measurements and validate our simulation assumptions.
[…]